11. Let Y be a random variable with pdf fy(y) = 6y5, o sys1. Determine the pdf of W = VY. (Hint: first compute the CDF.)
Answer #1

11)

CDF of Y is given by:

F(y) = P(Y <y) = f(y)dy = 6 ()dy = [y]% = y – 06 = y

CDF of W is given by:

F(W) = P(W<w) = P(VY <w) = P(Y < 3)

=(w^3)^6=w^{18}

So, the PDF of W is given by:

f(w) = f(W) = 18w17,0 <w<1